Martin Incorporated provided the following information regarding its only​ product: Sale price per unit ​$50.00 Direct materials used $ 165 comma 000 $165,000 Direct labor incurred $ 189 comma 000 $189,000 Variable manufacturing overhead $ 121 comma 000 $121,000 Variable selling and administrative expenses $ 72 comma 000 $72,000 Fixed manufacturing overhead ​$65,000 Fixed selling and administrative expenses ​$12,000 Units produced and sold 24 comma 000 24,000 Assume no beginning inventory Assuming there is excess​ capacity, what would be the effect on operating income of accepting a special order for 5 comma 500 5,500 units at a sale price of $ 42 $42 per​ product? (NOTE: Assume regular sales are not affected by the special order. Round any intermediary calculations to the nearest​ cent.)
Answer #1
Variable costs per unit:
Direct materials 6.88 =165000/24000
Direct labor 7.88 =189000/24000
Variable manufacturing overhead 5.04 =165000/24000
Variable selling and administrative expenses 3.00 =72000/24000
Total Variable costs per unit 22.80
Incremental revenue 231000 =5500*42
Less: Incremental costs 125400 =5500*22.80
Operating income 105600 increase